Hello, could someone help me with this problem? I'd really appreciate a step-by-step explanation. "Federico worked hard, but could only reduce his debt by 3 1/3. If this was only 2/9 of what he wanted to reduce, how much did he want to reduce it by?"
Thank you in advance.

Answer by Alan3354(69443)

"Federico worked hard, but could only reduce his debt by 3 1/3. If this was only 2/9 of what he wanted to reduce, how much did he want to reduce it by?"
----------
3 1/3 is 2/9 of Amount
(2/9)*Amt = 3 1/3 = 10/3
Multiply by 9
2*Amt = 30
Amt = 15